4:57 PM 72% 12ヤ13 3.7A V2- Vi-21V 11 = 4.2A V3= V4= 14 =.5A R4 = V5= V7 = 17-10A R5-6.12 Vt 240V V6 Rt- 18 OA 8 1.5Ω R6 =252 V9 = V12-2V R9 = R12= V10= 110 = R10 1552 V13-2V 13-3.5A R13-,571 RI 1 -4.5Ω can anyone help me fill in all the blanks? I keep getting stuck!
Answer #1

Solution

II2 = 6.5 A

R12 = 0.31 ohms

R13 = 0.57 ohms

I9 = 6.34 A, I10 = 0.85A , I11 = 2.82 A

V9 = V10 = V11= 12.7 V

V5 = V6 = 8.736 V

I5 = 8.736 /6.1 = 1.432 A and I6 = 4.368 A

V1 + (V2+3+4) and V5+6 are equal since they are parallel branches.

V2=V3 = V4   since they are parallel

V2 = 4.368 - 2.1 = 2.268 V

I2 = V2/R2 = 2.268 / 12 = 0.189 A

R4 = V4 I4 = 2.268 × 0.5 = 1.134 Ohms

I3 = 4.2 - ( 0.189+0.5) = 3.51 A

R3 = 2.268 / 3.51 = 0.646 Ohms

Total resistance   = 240/ 10 = 24 ohms

using the above values v7 and R7 can be determined.